C2H4S; experimental molar mass


1
Expert's answer
2021-03-05T03:30:59-0500

C2H4S

Carbon = 2×12 = 24

Hydrogen = 4×1 = 4

Sulphur = 1× 32 = 32

experimental molar mass= 24 + 4 + 32 = 60